# Helios: A Foundational Language Model for Smart Energy Knowledge Reasoning and Application

### v1

@@ 3.2.1 Universal Human Instruction Comprehension Data
In this stage, we have carefully selected six highly-recognized and high-quality open-source general-purpose supervised datasets: Alpaca-cleaned Taori et al. (2023b), Dolly-15K Conover et al. (2023b), Natural-Instructions Muennighoff (2022); Mishra et al. (2022); Wang et al. (2022a), python_code_25k FLOCK4H (2023), OpenR1-Math-220k lewtun and Face (2025), and Toolbench Qin et al. (2023). These datasets cover universal instruction understanding, mathematical reasoning, code enhancement, and tool utilization domains to improve Heilos’s foundational capabilities and domain application potential. For detailed information, please refer to supplementary material Section C.
@@ 3.2.2 Domain-specific Task Adaptation data: EnerInstruct
Seed Data Collection. In this study, we engaged 10 senior experts in the smart-energy domain to manually construct sample pairs for eleven downstream tasks: Fact Verification (FV), Reasoning (Res), Named Entity Recognition (NER), Summarization (Sum), Word Semantics (WS), Question and Answers (Q&A), Text Classification (TC), Explanation (Exp), Energy System Modeling (ESM), Single‑Choice (S-C) and Multiple-Choice (M-C). Which across fourteen sub-fields: clean energy, cogeneration, combined cooling–heating–and–power, distributed energy, energy hub, energy management system, energy optimization, energy storage, energy transition, integrated energy, load forecasting, smart energy, smart grid, and virtual power plant. The resulting seed dataset, covers all fourteen sub-fields and ten task categories, comprising 10000 data samples.
@@
Tasks Records Dataset Quality Optimization Total (Cleaned))
@@
Dataset Quality Optimization. During dataset construction, domain‑expert agents generated a large number of highly specialised samples for the various tasks. Nevertheless, the stochastic nature of sampling, the structural constraints imposed by the context‑window length, and the potential hallucinations produced by large language models can all cause fluctuations in sample quality and completeness. Extensive empirical work demonstrates that dataset size governs coverage and diversity, whereas sample quality determines the attainable upper bound on performance; the two must be carefully balanced. Suppose low‑quality samples containing redundancy, noise, or inconsistent annotations are used for training. They will dilute the informative signal, amplify systemic bias, and ultimately erode the model’s ability to follow instructions. To this end, we constructed a Check-Agent based on OpenAI o1, categorized by task, scoring each sample across the four dimensions of accuracy, completeness, relevance, and usefulness (out of 10), and providing reasons.
Samples that reach or exceed the threshold are retained, whereas those that do not are forwarded to an independent Optimisation-Agent. Guided by the evaluation feedback, this agent performs automatic remediation—correcting errors, supplementing and enriching content, or conducting deeper analysis as necessary. The revised sample is then returned to the Check-Agent for re‑evaluation. This “scoring–optimisation–re‑scoring” loop may iterate up to ten times: if a sample passes within the allotted rounds, it is admitted to the training corpus; if it fails all ten rounds, it is deemed irreparable and permanently discarded. Check‑Agent and the Optimisation-Agent collaboratively optimise the data workflow, as shown in Fig 3. This procedure ultimately yields dataset H‑EnerInstruct.
Figure 3: Dataset Quality Optimization workflow example. (a) Text before processing; (b) the Check‑Agent scores the text quality and provides optimization suggestions; (C) the Optimisation‑Agent generates the optimized text based on those suggestions. We mark the differences in Red.
Expert Manual Validation. Finally, a panel of 12 domain experts rigorously examined each task sample in H‑EnerInstruct (sampling 100–200 entries per task, proportional to that task’s size). Tasks that did not meet the required standard were flagged, and the experts issued uniform revision guidelines that were then refined by OpenAI o1 to ensure the dataset’s reliability. The optimized data were merged with the seed dataset to produce the final DS-TA phase dataset, EnerInstruct (Table 2). The statistics on expert optimization iterations are reported in Supplementary Section G.
@@ 3.3 RLHF Data: EnerReinforce
After large‑scale pre‑training and supervised instruction fine‑tuning, Helios can already address a wide range of tasks in the energy domain. Nevertheless, these stages seldom make human values or preferences explicit, so the resulting models may acquire generation patterns that diverge from human expectations. To align Helios more effectively with human preferences, we adopt a targeted, two‑stage approach consisting of reward model training followed by rejection sampling fine‑tuning. To this end, we have constructed EnerReinforce, which includes:
@@ 3.4 Evaluation on Expertise in Smart Energy: EnerBench
To systematically assess the problem-solving capabilities of LLMs on scientific questions in smart-energy research, we developed EnerBench, whose item-generation workflow adheres to a dual-track paradigm of Public-bank retrieval and Expert-directed authoring:
@@
2) Expert-directed Authoring: Five senior scholars in the smart-energy domain were commissioned to craft additional, high-quality items for every task within each sub-discipline, thereby augmenting the benchmark’s novelty and difficulty.
In its final form, EnerBench comprises 976 objective questions (S-C, M-C, and FV) and 625 subjective questions (Q&A, Exp, and ESM).
@@ 4.1 Pre-training
During the pre-training stage, we employ the Qwen-2.5 7B model Yang et al. (2024) (7.62 B trainable parameters) as the initialization weights for Helios. A single-epoch training is subsequently conducted on a domain-specific corpus of approximately 3 billion tokens in the smart-energy domain (22532 gradient update steps); the training hardware configuration consists of four NVIDIA A100-SXM 80 GB GPUs, with a total training time of 87 hours. The principal hyper-parameter settings are as follows: a peak learning rate of 3e-5, global batch size of 64, and a corresponding micro-batch size of 2.
@@ 4.3 RLHF
Reward Model Training. We employ a pairwise ranking loss to train the reward model, enabling it to distinguish between responses of varying quality:
@@ 5 Evaluation and Results
Object Tasks in EnerBench. For objective tasks, performance is evaluated using accuracy. Specifically, for multiple-choice items, the scoring rubric is: full credit is awarded only when all correct options are selected; partial credit is granted when some correct options are omitted; and no credit is given if any incorrect option is chosen. Helios attains an average accuracy of 79.09% in answering object questions, markedly outperforming models of comparable size such as Qwen3-8B-Instruct (41.87%) and LLama3-8B-Instruct (54.93%), and reaching a level comparable to GPT-4 with approximately 220 billion parameters. This indicates that the model successfully acquired intelligent-energy domain knowledge during further pre-training.
@@ 6 Conclusion
In this study, we introduce Helios, the first LLMs explicitly developed for the smart-energy domain, capable of addressing diverse tasks. We introduce EnerSys, a comprehensive multi-agent pipeline that furnishes Helios with high-quality data, producing EnerBase,EnerInstruct, and EnerReinforce. We also release Benchmark, the domain’s first evaluation suite, enabling systematic appraisal of LLMs on smart-energy tasks. Experiments show that, Helios offers significant gains in domain knowledge and task performance.

### v2

@@ 3.2.1 Universal Human Instruction Comprehension Data
In this stage, we have carefully selected six highly-recognized and high-quality open-source general-purpose supervised datasets: Alpaca-cleaned Taori et al. (2023b), Dolly-15K Conover et al. (2023b), Natural-Instructions Muennighoff (2022); Mishra et al. (2022); Wang et al. (2022a), python_code_25k FLOCK4H (2023), OpenR1-Math-220k R1 (2025), and Toolbench Qin et al. (2023). These datasets cover universal instruction understanding, mathematical reasoning, code enhancement, and tool utilization domains to improve Helios’s foundational capabilities and domain application potential. The data volume of each dataset in the UHIC phase is shown in Table 2.
@@ 3.2.2 Domain-specific Task Adaptation data: EnerInstruct
Seed Data Collection. In this study, we engaged 10 senior experts in the smart energy domain to manually construct sample pairs for eleven downstream tasks: Fact Verification (FV), Reasoning (Res), Named Entity Recognition (NER), Summarization (Sum), Word Semantics (WS), Question and Answers (Q&A), Text Classification (TC), Explanation (Exp), Energy System Modeling (ESM), Single‑Choice (S-C) and Multiple-Choice (M-C). Which across fourteen sub-fields: clean energy, cogeneration, combined cooling–heating–and–power, distributed energy, energy hub, energy management system, energy optimization, energy storage, energy transition, integrated energy, load forecasting, smart energy, smart grid, and virtual power plant. The resulting seed dataset, covers 14 sub-fields and 10 task categories, comprising 10 000 samples.
@@
Tasks Records Dataset Quality Optimization Total (Cleaned)
@@
Dataset Quality Optimization. During dataset construction, domain‑expert agents generated a large number of highly specialised samples for the various tasks. Nevertheless, the stochastic nature of sampling, the structural constraints imposed by the context‑window length, and the potential hallucinations produced by LLMs can all cause fluctuations in sample quality and completeness. Extensive empirical work demonstrates that dataset size governs coverage and diversity, whereas sample quality determines the attainable upper bound on performance; the two must be carefully balanced. Suppose low‑quality samples containing redundancy, noise, or inconsistent annotations are used for training. They will dilute the informative signal, amplify systemic bias, and ultimately erode the model’s ability to follow instructions. To this end, we constructed a Check-Agent based on OpenAI o1, categorized by task, scoring each sample across the four dimensions of accuracy, completeness, relevance, and usefulness (out of 10), and providing reasons.
Samples that reach or exceed the threshold are retained, whereas those that do not are forwarded to an independent Optimization-Agent. Guided by the evaluation feedback, this agent performs automatic remediation—correcting errors, supplementing and enriching content, or conducting deeper analysis as necessary. The revised sample is then returned to the Check-Agent for re‑evaluation. This “scoring–Optimization–re‑scoring” loop may iterate up to ten times: if a sample passes within the allotted rounds, it is admitted to the training corpus; if it fails all ten rounds, it is deemed irreparable and permanently discarded. Check‑Agent and the Optimization-Agent collaboratively optimise the data workflow, as shown in Fig 4. This procedure ultimately yields dataset H‑EnerInstruct.
Figure 4: Dataset Quality Optimization workflow example. (A) Text before processing; (B) the Check‑Agent scores the text quality and provides optimization suggestions; (C) the Optimization‑Agent generates the optimized text based on those suggestions. We mark the differences in Red.
Expert Manual Validation. Finally, a panel of 12 domain experts rigorously examined each task sample in H‑EnerInstruct (sampling 100–200 entries per task, proportional to that task’s size). Tasks that did not meet the required standard were flagged, and the experts issued uniform revision guidelines that were then refined by OpenAI o1 to ensure the dataset’s reliability. The optimized data were merged with the seed dataset to produce the final DS-TA phase dataset, EnerInstruct (Table 3). The statistics on expert optimization iterations are reported in Supplementary Section C.
@@ 3.3 RLHF Data: EnerReinforce
After large‑scale pre‑training and supervised instruction fine‑tuning, Helios can already address a wide range of tasks in the energy domain. Nevertheless, these stages seldom make human values or preferences explicit, so the resulting models may acquire generation patterns that diverge from human expectations. To align Helios more effectively with human preferences, we adopt a targeted, two‑stage approach consisting of reward model training followed by rejection sampling fine‑tuning, and constructed EnerReinforce, which includes:
@@ 3.4 Evaluation on Expertise in Smart Energy: EnerBench
To systematically assess the problem-solving capabilities of LLMs on scientific questions in smart energy research, we developed EnerBench, whose item-generation workflow adheres to a dual-track paradigm of Public-bank retrieval and Expert-directed authoring:
@@
2) Expert-directed Authoring: Five senior scholars in the smart energy domain were commissioned to craft additional, high-quality items for every task within each sub-discipline, thereby augmenting the benchmark’s novelty and difficulty.
In its final form, EnerBench comprises 887 objective questions (S-C, M-C, and FV) and 625 subjective questions (Q&A, Exp, and ESM). The detailed distribution of questions across sub-disciplines is provided in Table 4.
Table 4: The statistics of EnerBench.
Question Type Task Prompts
Objective task S-C 405
M-C 254
FV 228
Subjective tasks Q&A 196
Exp 249
ESM 180
@@ 4.1 Pre-training
During the pre-training stage, we employ the Qwen-2.5 7B model Yang et al. (2024) (7.62B trainable parameters) as the initialization weights for Helios. A single-epoch training is subsequently conducted on a domain-specific corpus of approximately 3 billion tokens in the smart energy domain (22532 gradient update steps); the training hardware configuration consists of four NVIDIA A100-SXM 80 GB GPUs, with a total training time of 87 hours. The principal hyper-parameter settings are as follows: a peak learning rate of 3e-5, global batch size of 64, and a corresponding micro-batch size of 2.
@@ 4.3 RLHF
Reward Model Training. We employ a pairwise ranking loss to train reward model, enabling it to distinguish between responses of varying quality:
@@ 5 Evaluation and Results
Objective Tasks in EnerBench. For objective tasks, performance is evaluated using accuracy. Specifically, for multiple-choice items, the scoring rubric is: full credit is awarded only when all correct options are selected; partial credit is granted when some correct options are omitted; and no credit is given if any incorrect option is chosen. Helios attains an average accuracy of 79.09% in answering object questions, markedly outperforming models of comparable size such as Qwen3-8B-Instruct (41.87%) and Llama3-8B-Instruct (54.93%), and reaching a level comparable to GPT-4 with approximately 220 billion parameters. This indicates that the model successfully acquired intelligent-energy domain knowledge during further pre-training.
@@
Figure 5: Case analysis of modeling tasks in the smart energy domain.
Exploring the Potential of Helios. We attempt to address a practical modelling and Optimization task in the smart energy domain using Helios. In this example, our requirement is: How to apply rainwater energy recovery in urban sunken interchange drainage pump stations? and to provide the implementation code (Fig. 5). It can be observed that Helios is able to effectively invoke domain-specific packages for intelligent energy (oemof and message_ix) to accomplish the modelling task, whereas Llama3-8B can only call numpy to perform purely numerical computations, which deviates substantially from the task requirements and lacks practical relevance to the energy domain.
@@ 6 Conclusion
In this study, we introduce Helios, the first LLM explicitly developed for the smart energy domain, capable of addressing diverse tasks. We introduce EnerSys, a comprehensive multi-agent pipeline that furnishes Helios with high-quality data, producing EnerBase, EnerInstruct, and EnerReinforce. We also release Benchmark, the domain’s first evaluation suite, enabling systematic appraisal of LLMs on smart energy tasks. Experiments show that, Helios offers significant gains in domain knowledge and task performance.
Limitations
Although Helios has demonstrated excellent capabilities in knowledge integration and automatic code generation within the smart energy domain, its role is consistently positioned as an "intelligent reference assistant" rather than an autonomous decision-making engine. In high-risk tasks such as power system modeling, dispatch, and safety assessment, Helios only outputs code drafts and inferential suggestions for review by engineers. Direct deployment without professional review could lead to significant economic losses or even physical risks due to potential model assumption biases, numerical instability, or omission of boundary conditions. Consequently, the model’s outputs do not constitute an engineering guarantee. The final decision-making responsibility must be borne by the user and their affiliated institution; when results are uncertain or contradict engineering experience, it is essential to revert to traditional manual calculation and simulation for verification.
Concerning hallucinations in Helios, they mainly manifest as linguistic repetition, instruction misunderstanding, conceptual confusion, and structural errors. For instance, in factual judgment tasks, the model occasionally misinterprets the task as question-answering, a problem that is significantly mitigated by explicitly appending "Please output True/False directly" to the prompt. Conceptual confusion stems from the interdisciplinary, fragmented, and rapidly evolving nature of smart energy knowledge; experiments show its occurrence rate remains within an acceptable range. Linguistic repetition and structural hallucinations are largely associated with the base model, Qwen-2.5-7B, and are difficult to eliminate completely through domain-specific fine-tuning alone; thus, they are not a primary focus of this paper.In summary, Helios has the aforementioned limitations regarding ethics, risks, and deployment, and should be applied cautiously within a strict framework of human-computer collaboration and safety governance.
Ethical considerations
Regarding ethics and bias, Helios is primarily trained on high-quality corpora such as academic papers and monographs, and its instruction data has undergone rigorous cleaning, resulting in minimal potential for ethical or bias issues.
Acknowledgments
This work is supported by the National Natural Science Foundation of China (Grant No.52576234) and the National Key R & D Program of China (Grant No.2023YFE0108600). This work is also supported by the "Pioneer" and "Leading Goose" R & D Program of Zhejiang (Grant No. 2025C02237)